The question is: log x^3 - 2log 1/x - 1= 0
What I tried:
3log x- (-2log x) - 1= 0
3log x + 2 log x - 1= 0
5 log x- 1 = 0
5 log x = 1
log x = 1/5
x= 10^1/5
or
3log x- (-2log x) - 1= 0
3log x + 2 log x - 1= 0
5 log x- 1 = 0
5 log x = 1
log x^5 = 1
x^1=5
x=5
I'm still confused with this kind of questions. Thanks in advance ^0^!

Answer by nerdybill(7384)

log x^3 - 2log 1/x - 1= 0
log x^3 - 2log 1/x = 1
log x^3 - log (1/x)^2 = 1
log x^3 - log (1/x^2) = 1
log x^3 - log (x^-2) = 1
log x^3/(x^-2) = 1
log x^3x^2 = 1
log x^5 = 1
x^5 = 10^1
x^5 = 10
x = 10^(1/5)
x = 1.5849